OpenStudy (anonymous):

what is 2 to the 3rd power divided by 4 minus 2

OpenStudy (anonymous):

0

OpenStudy (anonymous):

2^3 = 2 x 2 x 2 = 8 8/4 = 2 2-2 =0

OpenStudy (anonymous):

zero